A steel wire having a mass of 6.30 g and a length of 1.20 m is fixed at both ends and has a tension of 955 N. (a) Find the speed of transverse waves on the wire. 1 405 Incorrect: Your answer is incorrect. m/s (b) Find the wavelength of the fundamental. 2 m (c) Find the frequency of the fundamental. 3 Hz (d) Find the frequency of the second harmonic. 4 Hz (e) Find the frequency of the third harmonic. 5 Hz Find the speed of transverse waves on the wire from the tension and linear mass density. Find the wavelengths of any harmonic from the standing wave condition of waves on a string fixed at both ends. For a wire fixed at both ends, the higher harmonics are integer multiples of the first harmonic (fundamental).
Answer #1

0 M = 6.39 L = 1.2 m T= 955N (a) v= T/m T (M/L) TL M asso 1.2 6.3x10 3 r = 426.5 nis fundamental ware length 2,2 al 2x 1.2 2.fundamental 2 = N frequency = 426,5 2.4 177.71 H2 d on a no » harmonic n-2 2 27, 2 2 2 x 177.71 2 = 355.419 Hz za 32, 3x 177.